### v1

# The Three Regimes of Offline-to-Online Reinforcement Learning

Abstract
Offline-to-online reinforcement learning (RL) has emerged as a practical paradigm that leverages offline datasets for pretraining and online interactions for fine-tuning. However, its empirical behavior is highly inconsistent: design choices of online-fine tuning that work well in one setting can fail completely in another. We propose a stability–plasticity principle that can explain this inconsistency: we should preserve the knowledge of pretrained policy or offline dataset during online fine-tuning, whichever is better, while maintaining sufficient plasticity. This perspective identifies three regimes of online fine-tuning, each requiring distinct stability properties. We validate this framework through a large-scale empirical study, finding that the results strongly align with its predictions in 45 of 63 cases. This work provides a principled framework for guiding design choices in offline-to-online RL based on the relative performance of the offline dataset and the pretrained policy.
1 Introduction
Reinforcement learning (RL) has achieved impressive successes in a variety of domains (Mnih et al., 2015; Silver et al., 2017; Degrave et al., 2022), but its reliance on large amounts of online interaction often makes direct application to real-world problems challenging. To address this challenge, recent research has turned to leveraging pre-collected datasets for offline learning through offline RL (Levine et al., 2020) or imitation learning (Osa et al., 2018). This paradigm reduces the need for costly or unsafe online exploration by providing a strong initial policy trained entirely from offline data. However, policies trained purely offline are often suboptimal and fail to generalize to states outside the dataset’s support, making online fine-tuning essential. Offline-to-online RL addresses this issue by first pretraining an agent on an offline dataset and then fine-tuning it with additional online interactions to further improve performance.
Figure 1: Comparison between WSRL (pretrained policy only) and RLPD (offline dataset only) on two representative offline-to-online RL tasks. All learning curves are shown as mean $\pm$ 95% CI.
While offline-to-online RL has led to promising results (Nair et al., 2020; Lee et al., 2022), the online RL fine-tuning has shown highly inconsistent empirical behavior. Design choices that work well in one setting can fail completely in another. For example, as shown in Figure 1, on D4RL tasks (Fu et al., 2020) such as antmaze-large-play-v2, Warm-Start RL (WSRL) (Zhou et al., 2024), which relies on the pretrained policy and discards the offline dataset during online fine-tuning, substantially outperforms RL with Prior Data (RLPD) (Ball et al., 2023), which uses only the offline dataset. In contrast, on D4RL tasks such as relocate-binary-v0, the opposite pattern emerges, with RLPD outperforming WSRL by a wide margin.
These seemingly inconsistent outcomes raise one fundamental question:
What underlying factors cause design choices to succeed in some settings but fail in others?
To answer this question, we propose a stability–plasticity principle for offline-to-online RL that can explain these seemingly inconsistent outcomes, which is inspired by prior research on plasticity and stability in neuroscience (McClelland et al., 1995) and machine learning  (Kirkpatrick et al., 2017; Wolczyk et al., 2024; Dohare et al., 2024). This principle states that effective fine-tuning requires a careful balance between stability and plasticity. Stability refers to the preservation of useful prior knowledge, ensuring that competencies acquired during pretraining are not substantially degraded. Plasticity, in contrast, denotes the capacity of the model to adapt flexibly and efficiently to new data. Furthermore, we identify two distinct forms of stability in offline-to-online RL: stability around the pretrained policy $\pi_{0}$,which emphasizes preserving knowledge explicitly encoded in the policy parameters, and stability around the offline dataset $\mathcal{D}$, which emphasizes retaining knowledge implicitly encoded in offline data. As stability and plasticity are inherently in trade-off, this distinction indicates that fine-tuning is more efficient when stability is enhanced with respect to the stronger source of prior knowledge, whether it be the pretrained policy or the offline dataset.
Building on this insight, we propose a taxonomy of three regimes for offline-to-online RL, each capturing a distinct relationship between the pretrained policy and the offline dataset. This framework provides actionable guidance for practitioners. As shown in Figure 2, the three regimes are defined based on which source of prior knowledge is stronger, either the pretrained policy or the offline dataset. Moreover, different fine-tuning methods can be systematically categorized according to whether they enhance stability around $\pi_{0}$, stability around $\mathcal{D}$, or plasticity, and each category can be positioned within these regimes. By first determining the regime, one can select or design fine-tuning strategies that match its stability-plasticity requirements. This yields two practical benefits: it helps choose the most suitable method for each setting rather than applying a single uniform state-of-the-art algorithm, and it narrows the search space by indicating whether one should focus on leveraging the pretrained policy or on exploiting the offline dataset, thereby reducing unnecessary trial-and-error.
To validate this framework, we conduct a large-scale empirical study that covers 21 dataset-task compositions across four domains in D4RL: Mujoco locomotion, AntMaze, Adroit, and Kitchen. Combined with three representative pretraining algorithms, this results in 63 experimental settings. The observed behaviors align closely with the predictions of our framework, supporting its utility as a principled basis for guiding design choices of fine-tuning in offline-to-online RL.
Contributions of this paper can be summarized as:
  • We propose a stability–plasticity principle for offline-to-online RL.
  • We develop a taxonomy of three regimes based on the relative performance of $\pi_{0}$ and $\mathcal{D}$.
  • We conduct an extensive empirical study across 63 settings to validate and analyze this framework.
2 Preliminary: Offline-to-online RL
Figure 2: Overview of the three regimes in offline-to-online RL, defined based on the relative performance of the pretrained policy $J(\pi_{0})$ and the offline dataset $J(\pi_{\mathcal{D}})$. For each regime, our framework indicates which property is most needed during fine-tuning. The boxes at the right show representative design choices that implement these enhancing stability or plasticity.
Consider an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$, where the performance of a policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ is measured by its expected discounted return: $J(\pi)=\mathbb{E}_{\pi,\mathcal{M}}[\sum_{t}\gamma^{t}r_{t}]$. Offline-to-online RL begins by pretraining the agent on an offline dataset $\mathcal{D}$, which is collected from $\mathcal{M}$ under an unknown behavior policy (or mixture of policies), using an offline RL algorithm $A_{\text{off}}$. This yields an offline pretrained agent whose policy is given by $\pi_{0}=A_{\text{off}}(\mathcal{D})$, The fine-tuning step consists in using an algorithm $A_{\text{on}}$ using $\pi_{0}$ to interact with $\mathcal{M}$ in order to derive a final policy $\pi_{N}$:
$\pi_{N}=A_{\text{on}}\left(\mathcal{M},\mathcal{D},\pi_{0}\right).$ (1)
The offline dataset $\mathcal{D}$ may be reused (Nakamoto et al., 2023) or discarded (Zhou et al., 2024) during fine-tuning. The objective of offline-to-online RL is to co-design $(A_{\text{off}},A_{\text{on}})$ such that the final policy $\pi_{N}$ maximizes $J(\pi_{N})$. In this work, we focus on understanding and improving the online RL fine-tuning component $A_{\text{on}}$ by fixing the offline pretraining component $A_{\text{off}}$.
3 A Stability–Plasticity Principle on Knowledge Acquisition
This section introduces a formal framework for reasoning about fine-tuning in offline-to-online RL. Our goal is to characterize when and how online training leads to improvements over the offline initialization—without degrading what was already learned. We define two complementary properties of online fine-tuning: stability, the ability to preserve previously acquired performance, and plasticity, the capacity to improve further. These are grounded in the notion of a knowledge level, understood as the expected return encoded either in the dataset or in the pretrained policy. We show that final performance admits a decomposition into three terms—prior knowledge, stability, and plasticity. We leverage this perspective provides both theoretical insight and practical guidance.
3.1 Quantifying Prior Knowledge
We distinguish two sources of prior knowledge available before online fine-tuning. The first is the knowledge encoded in the dataset, captured by the performance of the behavior policy that generated it. The second is the knowledge embodied in the pretrained policy, obtained by running an offline RL algorithm on the dataset. Both can serve as baselines for measuring stability and plasticity.
Knowledge from the dataset: $J(\pi_{\mathcal{D}})$.
Let $\mathcal{D}$ be the offline dataset and $\pi_{\mathcal{D}}$ be an abstract behavior policy representing the data-generating process. While $\mathcal{D}$ may be collected from a mixture of policies, $\pi_{\mathcal{D}}$ serves as a convenient abstraction. Its performance can be estimated by the average return:
$J(\pi_{\mathcal{D}})\approx\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}r_{i,t}.$
This measure provides a scalar summary of the return level encoded in the dataset. It does not capture coverage or diversity, but it defines a baseline that fine-tuning should not fall below.
Knowledge from the pretrained policy: $J(\pi_{0})$.
Let $\pi_{0}$ be the policy produced by applying an offline RL algorithm $A_{\text{off}}$ to $\mathcal{D}$. Its performance, $J(\pi_{0})$, reflects not only the quality of the dataset but also the inductive biases of $A_{\text{off}}$ and the difficulty of the MDP $\mathcal{M}$. This policy can provide a strong initialization for online fine-tuning, though it does not necessarily dominate the dataset baseline. We therefore consider both $J(\pi_{\mathcal{D}})$ and $J(\pi_{0})$ jointly as candidate sources of prior knowledge.
3.2 Knowledge Decomposition and the Three Regimes of Offline-to-online RL
Online fine-tuning produces a sequence of policies $\{\pi_{n}\}_{n=0}^{N}$ with corresponding performances $\{J(\pi_{n})\}_{n=0}^{N}$. Our goal is to understand how these trajectories of performance can be expressed in terms of the prior knowledge identified above and the two complementary properties of stability and plasticity. This leads to a decomposition of final performance that makes explicit what is preserved from offline training and what is gained during online interaction.
Stability with regard to a knowledge level.
We define the stability of an online RL training process with respect to a knowledge level $l$, as the ability to retain the relative performance:
$\text{Stability}(l)=\min\left(\min_{0\leq n\leq N}J(\pi_{n})-l,0\right).$ (2)
This metric captures the worst-case performance drop during online fine-tuning relative to a given baseline $l$. A stability score of zero indicates that performance was never degraded below $l$; a negative score reflects how much was lost.
In our setting, the appropriate reference level is the best performance available from the offline pretraining phase, either from the dataset or from the pretrained policy: {myframe}
$J^{*}_{\text{off}}=\max\left(J(\pi_{0}),\;J(\pi_{\mathcal{D}})\right).$
We refer to this as the offline performance baseline, and we define stability with respect to it as:
$\text{Stability}(J^{*}_{\text{off}})=\min_{0\leq n\leq N}J(\pi_{n})-J^{*}_{\text{off}}\leq 0.$ (3)
This condition ensures that the fine-tuning process does not degrade below the strongest available offline signal.
Plasticity.
We define the plasticity of an online RL algorithm as the ability to acquire new knowledge:
$\text{Plasticity}=\max_{0\leq i\leq N}J(\pi_{i})-\min_{0\leq j\leq N}J(\pi_{j})\geq 0.$ (4)
This metric measures the extent to which the performance can improve during fine-tuning, relative to its lowest observed value. A larger value indicates stronger capacity to learn from new data.
Our definition enables such comparison by relating both concepts through a knowledge decomposition:{myframe}
$\underbrace{\max_{0\leq n\leq N}J(\pi_{n})}_{\text{Final knowledge}}=\underbrace{J^{*}_{\text{off}}}_{\text{Prior knowledge}}+\underbrace{\text{Stability}(J^{*}_{\text{off}})}_{\text{Degradation on prior knowledge}\leq 0}+\underbrace{\text{Plasticity}}_{\text{Online knowledge}\geq 0}.$
This equation states that the final knowledge an agent acquires after offline pretraining and online fine-tuning is the outcome of three interacting components. The first term is the prior knowledge provided by the offline phase, either through the dataset or the pretrained policy, whichever is stronger. The second term measures stability, which records whether this prior knowledge is preserved or degraded during fine-tuning; it is always non-positive since performance can at best be maintained but not exceeded by this term. The third term captures plasticity, the additional knowledge acquired through online interaction, which is non-negative by definition.
Viewed this way, the outcome of offline-to-online RL is neither purely a continuation of offline learning nor purely the result of online finetuning. It is the byproduct of what was preserved from the offline baseline and what was added through adaptation, minus any performance lost in the process.
The three regimes of offline-to-online RL.
With prior knowledge fixed, the objective of online fine-tuning is to improve final performance by maintaining stability with respect to $\max(J(\pi_{0}),J(\pi_{\mathcal{D}}))$ while ensuring sufficient plasticity. Based on this perspective, we identify three regimes for the online fine-tuning phase of offline-to-online RL: Superior: where $J(\pi_{0})>J(\pi_{\mathcal{D}})$; Comparable: where $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$; Inferior: where $J(\pi_{0})<J(\pi_{\mathcal{D}})$. These regimes are intended to reflect substantial differences in $J(\pi_{0})$ and $J(\pi_{\mathcal{D}})$, since small performance gaps may not be meaningful and therefore should not determine regime assignment.
This regime taxonomy provides a principled framework for reasoning about the stability–plasticity trade-off in offline-to-online RL. It clarifies which source of prior knowledge should anchor stability in a given setting, as shown in Figure 2. In the Superior Regime, stability relative to $\pi_{0}$ should be prioritized, meaning that online updates should avoid degrading below the performance already achieved by the pretrained policy. In the Inferior Regime, stability relative to $\mathcal{D}$ should be emphasized, meaning that fine-tuning should at least preserve the return level encoded in the dataset. In the Comparable Regime, both baselines provide similar performance, so preserving either suffices. By categorizing design choices according to the stability or plasticity they promote, our framework turns what previously appeared as a disconnected set of practices into a structured and predictable landscape, providing a principled foundation for future algorithm design.
4 Design Choices in Improving Stability or Plasticity
Building on the stability–plasticity principle and the regime taxonomy, we now examine concrete design choices that instantiate these principles. To study these factors systematically, we isolate and analyze representative modules, each of which targets one of the three directions: stability around $\pi_{0}$, stability around $\mathcal{D}$, or increased plasticity. Although many existing offline-to-online RL algorithms combine multiple components that promote these aspects simultaneously, such entanglement makes it difficult to attribute effects to a specific source. To obtain a clearer scientific understanding, we isolate and analyze representative modules individually.
Minimal baseline.
We begin with defining a naive online RL fine-tuning baseline that serves as the reference point for introducing additional components, which is intentionally minimalist. It applies a standard online RL algorithm initialized with an offline-pretrained agent, without any further modifications. This baseline anchors the analysis and makes the marginal effect of each added component interpretable.
4.1 Stability relative to the offline dataset $\mathcal{D}$
Design choices in this category promote stability by reusing the offline dataset during online fine-tuning. Incorporating $\mathcal{D}$ into the online learning process helps preserve the knowledge in the offline dataset and mitigates distribution shift between offline and online data.
A common strategy for incorporating offline data during fine-tuning is to reuse the offline dataset together with newly collected online transitions. One approach initializes the replay buffer with the entire offline dataset, after which new online experiences are appended as the agent interacts with the environment. In this case, the ratio of offline to online data is determined by the dataset size and gradually shifts toward online data as training progresses. An alternative approach maintains two separate replay buffers: one fixed buffer containing the offline dataset and another buffer for online experiences. During training, each batch is sampled from both buffers according to a to a specified offline data ratio $\alpha$. For instance, RLPD (Ball et al., 2023) uses $\alpha=0.5$, corresponding to a symmetric 50% offline and 50% online sampling ratio.
4.2 Stability with respect to the pretrained policy $\pi_{0}$
Design choices in this category focus on preserving and building upon the knowledge encoded in the pretrained policy $\pi_{0}$. The goal is to reduce the risk of catastrophic forgetting and ensure that fine-tuning does not erase useful behaviors learned during pretraining.
Online data warm-up.
Before applying gradient updates, the agent first collects a larger amount of online data ($K$ steps). This strategy reduces the mismatch between the pretraining distribution and the online data distribution, lowering the chance that early updates overwrite prior knowledge. It was introduced in WSRL (Zhou et al., 2024).
Offline RL regularization.
Fine-tuning can also reuse the same offline RL algorithm that produced the pretrained policy. This carries over the conservative regularization used in the offline phase, helping to control distribution shift and overestimation during online updates. Such regularization during fine-tuning has been widely adopted in prior work (Nair et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2023; Nakamoto et al., 2023).
4.3 Plasticity: Parameter Reset
The most direct way to increase plasticity is to periodically reset network parameters while retaining the accumulated training data. Randomly initialized networks tend to exhibit higher plasticity than pretrained ones (Nikishin et al., 2022), although this strategy may come at the cost of severe forgetting and performance degradation. In the context of offline-to-online RL, parameter reset can be interpreted as starting from any pretrained agent $\pi_{0}$ and then resetting its weights to a randomly initialized agent $\pi_{1}$. While this approach severely degrades initial performance (i.e., $J(\pi_{1})-J(\pi_{0})$ is highly negative), it can significantly enhances plasticity. RLPD (Ball et al., 2023) directly trains an online RL agent from random initialization, which can be viewed as pretraining followed by reset before fine-tuning.
5 Empirical Study
We test the validity of our three-regime framework through a large-scale empirical study, examining whether its regime-specific predictions align with observed outcomes. To connect the design modules described above with this framework, we group algorithms by the primary source of stability they emphasize. Methods that preserve knowledge from the pretrained policy $\pi_{0}$ are labeled $\bm{\pi_{0}}$-centric, while those that anchor stability to the offline dataset $\mathcal{D}$ are labeled $\bm{\mathcal{D}}$-centric. Approaches that combine elements of both are called mixed $\bm{\pi_{0}+\mathcal{D}}$ methods. Finally, the minimal baseline corresponds to maximum plasticity with no explicit stability mechanism.
We study a diverse set of benchmark tasks and pretraining dataset compositions, following the experimental protocols of prior work (Nakamoto et al., 2023; Zhou et al., 2024). Specifically, we include MuJoCo locomotion, AntMaze, Kitchen, and Adroit domains from D4RL (Fu et al., 2020), covering a total of 21 dataset-task compositions. All experiments are conducted with 10 random seeds to ensure statistical reliability.
Offline Pretraining Phase.
For the offline pretraining phase, we employ two representative offline RL algorithms, Calibrated Q-Learning (CalQL) (Nakamoto et al., 2023) and ReBRAC (Tarasov et al., 2023), as well as an additional deterministic behavior cloning (BC) (Schaal, 1996) pretraining, a commonly used imitation learning algorithm that serves as a simple alternative to offline RL. Notably, BC typically achieves performance comparable to that of the dataset itself. Combining the 21 dataset-task compositions with these 3 pretraining algorithms yields 63 experimental settings in total. Each setting is defined by a specific combination of pretraining algorithm, dataset, and task. In this work, since we focus on cases where the offline dataset and the task are from the same MDP, the pretraining algorithm and dataset are sufficient to uniquely specify a setting.
We use the regime classification introduced in Section 3 to organize our analysis and to interpret the outcomes of the fine-tuning methods. Each of the 63 experimental settings, defined by a unique combination of pretraining algorithm, dataset, and task, is assigned to one of the three regimes based on the relative performance of the pretrained policy and the offline dataset. Specifically, we conduct $t$-tests with a margin $\delta=0.05$ to assess whether the difference between $J(\pi_{0})$ and $J(\pi_{\mathcal{D}})$ is statistically significant. The margin $\delta$ is introduced for robustness, since $J(\pi_{\mathcal{D}})$ is approximated by the dataset average return and small gaps between $J(\pi_{0})$ and $J(\pi_{\mathcal{D}})$ may not be meaningful. It prevents over-interpreting numerical noise in regime assignment. The complete set of regime assignments is reported in Table 4 in the appendix.
Online Fine-Tuning Phase.
To ensure consistency between offline pretraining phase and online fine-tuning phase, we fine-tune each agent using the corresponding base algorithm. Specifically, we fine-tune CalQL-pretrained agents using SAC (Haarnoja et al., 2018) and ReBRAC-pretrained agents using TD3 (Fujimoto et al., 2018). For the deterministic BC pretraining, we use TD3 for fine-tuning to match its deterministic actor structure.
Since evaluating every possible design and their combinations is infeasible, and we have grouped them into four categories: the minimal baseline, $\pi_{0}$-centric methods, $\mathcal{D}$-centric methods, and mixed $\pi_{0}+\mathcal{D}$ methods. We then evaluate six representative methods spanning these four categories, which collectively capture the key design choices explored in prior offline-to-online RL literature.
  • Baseline: Fine-tunes the pretrained policy using an online RL algorithm with only online data.
  • $\bm{\pi_{0}}$-centric methods: Two variants of such methods are evaluated: the baseline with (i) online data warm-up (K=5, [^0] K=5{,}000 steps) and (ii) offline RL regularization using the pretraining coefficient.
  • $\bm{\mathcal{D}}$-centric methods: Two variants of such methods are evaluated: the baseline with (i) offline data replay, and (ii) with offline data replay and reset. Both variants use separate replay buffers with an offline data ratio of $\alpha=0.5$.
  • Mixed $\bm{\pi_{0}+\mathcal{D}}$ methods: the baseline combined with offline data replay and offline RL regularization, a combination widely adopted in prior work (Nair et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2023; Nakamoto et al., 2023).
In each setting, we focus and compare the strongest $\pi_{0}$-centric and $\mathcal{D}$-centric methods to better approximate the ideal performance achievable by each stability source, while minimizing confounding from implementation details and hyperparameter tuning. The following subsections present and analyze the fine-tuning results across the three regimes.
Table 1: Confusion matrix of fine-tuning results across the three regimes. Green cells: correct predictions (45/63); red cells: opposite mismatches (3/63); gray cells: adjacent mismatches (15/63). Overall, the framework achieves 71% correct predictions with only 5% opposite mismatches.
Pretraining Regime
Superior Comparable Inferior
Fine-tune
$\pi_{0}$-centric >> $\mathcal{D}$-centric 24 2 1
$\pi_{0}$-centric $\approx$ $\mathcal{D}$-centric 6 2 3
$\pi_{0}$-centric << $\mathcal{D}$-centric 2 4 19
5.1 Superior regime: $J(\pi_{0})>J(\pi_{\mathcal{D}})$
In the Superior regime, the pretrained policy $\pi_{0}$ achieves substantially higher performance than the behavior policy underlying the offline dataset, i.e., $J(\pi_{0})>J(\pi_{\mathcal{D}})$. In such cases, the offline dataset offers limited additional value, and the primary concern becomes preserving stability relative to the pretrained policy during online fine-tuning.
Figure 3: Representative fine-tuning results in the Superior regime: the first row and first two subplots in the second row are correct predictions, while the remaining two show an adjacent mismatch and an opposite mismatch. Markers on the curves indicate the better-performing variant within $\pi_{0}$-centric methods and within $\mathcal{D}$-centric methods.
Representative results are shown in Figure 3, with the complete set of results for all settings in this regime provided in the appendix (Figure 6). We perform $t$-tests between the strongest $\pi_{0}$-centric and $\mathcal{D}$-centric methods in each setting within this regime. The results indicate $\pi_{0}$-centric methods outperform $\mathcal{D}$-centric methods in 24 out of 32 settings (75%), while the remaining settings mostly show no statistically significant difference. Only two settings exhibit the opposite outcome, and in both cases the difference in average performance is negligible. One such case is shown in the rightmost plot of the second row in Figure 3. These statistics correspond to the Superior column of the confusion matrix in Table 1.
This aggregate outcome strongly supports our principle that, in the Superior regime, $\pi_{0}$-centric methods tend to be more effective than $\mathcal{D}$-centric methods. In other words: when the pretrained policy already outperforms the dataset (the superior case), methods that stick close to the pretrained policy work better than those that keep leaning on the offline dataset. While the prediction accuracy is not perfect, such discrepancies are anticipated given the influence of hyperparameters and implementation details. Importantly, the overall observed patterns remain consistent with our principle.
Beyond aggregate comparisons, the analysis of specific design choices highlights key trade-offs between stability and plasticity. Within the two variants of $\pi_{0}$-centric methods, online data warm-up achieves better performance in 27 out of 32 settings, reflecting its ability to preserve the pretrained policy’s knowledge while maintaining sufficient plasticity for fine-tuning. In contrast, offline RL regularization provides stronger stability, leading to substantially less performance degradation during the early fine-tuning phase. However, this strong stability comes at the cost of reduced plasticity, which limits long-term improvements during fine-tuning, making it better than online data warm-up in only 5 out of 32 settings, where the pretrained policy is already close to the optimal policy. The same applies to the combination of offline RL regularization with offline data replay, which exhibits the strongest stability among all methods considered. These contrasts highlight the importance of considering each setting and identifying the the method that best balances the underlying stability–plasticity trade-off.
Takeaway: In the Superior regime, where the pretrained policy $\pi_{0}$ substantially outperforms the offline dataset $\mathcal{D}$, $\pi_{0}$-centric methods are typically more effective than $\mathcal{D}$-centric methods. Stronger stability proves beneficial primarily when $\pi_{0}$ is already close to optimal.
Figure 4: Representative results in the Inferior regime: the first six results are correct predictions, while the remaining two show an adjacent mismatch and an opposite mismatch.
5.2 Inferior Regime: $J(\pi_{0})<J(\pi_{\mathcal{D}})$
In the Inferior regime, the pretrained policy $\pi_{0}$ performs substantially worse than the behavior policy underlying the offline dataset, i.e., $J(\pi_{0})<J(\pi_{\mathcal{D}})$. In this regime, the pretrained policy contributes substantially less useful knowledge than the offline dataset, making it crucial to retain and leverage the offline data during online fine-tuning.
Representative results are shown in Figure 4, with the complete set of results for all settings in this regime provided in the appendix (Figure 7). According to the results, $\mathcal{D}$-centric methods outperform $\pi_{0}$-centric methods in 19 out of 23 settings (83%), while the remaining settings mostly show no statistically significant difference. Only one setting exhibits the opposite outcome, which is illustrated in the rightmost plot of the second row in Figure 4. These statistics correspond to the Inferior column of the confusion matrix in Table 1. Taken together, these aggregate results show that in the Inferior regime, $\mathcal{D}$-centric methods tend to be more effective than $\pi_{0}$-centric methods, consistent with the prediction of our framework.
Notably, offline data replay with reset achieves better performance than offline data replay in 13 out of 24 settings, despite the fact that reset initially causes significant degradation. This indicates that in these cases the offline pretraining phase substantially reduces plasticity while offering limited useful knowledge, and resetting the parameters allows the agent to adapt and acquire new knowledge more effectively. Furthermore, combining offline RL regularization with offline data generally underperforms compared to $\mathcal{D}$-centric methods. Although this design leverages offline data during fine-tuning, which is essential in this regime, the excessive stability limits plasticity and thereby hinders further improvement.
Takeaway: In the Inferior regime, where the pretrained policy $\pi_{0}$ performs substantially worse than the offline dataset $\mathcal{D}$, $\mathcal{D}$-centric methods typically provide more effective fine-tuning than $\pi_{0}$-centric methods. Parameter reset can also be beneficial, particularly when the pretrained agent performs very poorly, as it restores plasticity and allows the agent to adapt more effectively.
5.3 Comparable Regime: $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$
In the Comparable regime, the pretrained policy and the dataset behavior policy achieve similar performance, i.e., $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$. Representative results are shown in Figure 5, with the complete set of results provided in the appendix (Figure 8).
Our framework predicts that $\pi_{0}$-centric and $\mathcal{D}$-centric methods should yield comparable outcomes once fully optimized. Empirically, only 2 out of 8 settings are statistically indistinguishable under $t$-tests. This seems at odds with the prediction. However, closer inspection shows that the differences are minor: in 6 of 8 settings the mean gap between categories is less than 0.1. These small gaps indicate that both anchors provide similar prior knowledge, exactly as the framework suggests.
Why, then, do mismatches arise at all? The key is that effect sizes in this regime are small by construction. When $\pi_{0}$ and $\mathcal{D}$ are nearly tied, outcomes become highly sensitive to hyperparameters, initialization, and other implementation details. In our study, we fixed a limited set of representative variants and hyperparameters across all settings to avoid over-tuning. This conservative design choice helps comparability but can also tip results in such close cases.
Figure 5: Representative fine-tuning results for the Comparable regime. The first two subplots illustrate cases consistent with our framework’s predictions, while the latter two show mismatches with only small mean differences.
Takeaway: In the Comparable regime, where the pretrained policy $\pi_{0}$ and the offline dataset $\mathcal{D}$ exhibit similar performance, $\pi_{0}$-centric and $\mathcal{D}$-centric methods should in principle yield comparable fine-tuning outcomes when fully optimized, though in practice their relative performance is often sensitive to implementation details.
6 Conclusion
This paper introduced the stability–plasticity principle as a way to reconcile the puzzling variability of offline-to-online RL. We showed that the key determinant of fine-tuning success is which source of prior knowledge (pretrained policy or offline dataset—serves as the stronger anchor). From this observation we derived a taxonomy of three regimes, each dictating where stability should be enforced and how plasticity should be managed. The value of this framework is twofold. First, it provides a clear explanation for the conflicting empirical evidence in the literature: design choices that seem inconsistent across benchmarks in fact reflect different underlying regimes. Second, it offers actionable guidance for practitioners. By identifying the regime of a given setting, one can select methods that align with its stability–plasticity requirements, reducing reliance on trial-and-error.
Our regime taxonomy provides an efficient lens for understanding offline-to-online RL by compressing complex phenomena into a small number of discrete categories. Such taxonomies have been highly successful in both the natural sciences and machine learning, for example in imitation learning, where regimes have been proposed based on density ratios (Spencer et al., 2021) or dataset size (Belkhale et al., 2023). At the same time, we recognize that discretizing behavior into three regimes is a simplification. Real systems often lie along a continuum, and other dataset characteristics—such as coverage—play an important role but remain difficult to capture consistently. Extending the framework to incorporate such dimensions is an important direction for future work.
Our work also connects offline-to-online RL to a broader body of research on the stability–plasticity dilemma in deep learning and neuroscience. Recent studies in continual and online deep RL have focused on characterizing forgetting and plasticity, but applications to the offline-to-online transition have been limited. We show that stability–plasticity is not only a useful lens for analyzing this setting, but also a source of practical guidance: it predicts which design choices are effective in which regimes, moving beyond coining terminology to actionable prescriptions.
Appendix A Related Work
Offline-to-online RL.
Offline-to-online RL seeks to combine the strengths of offline pretraining with the adaptability of online fine-tuning. Early approaches focused on extending offline RL regularization methods into the online regime, constraining fine-tuning updates to remain close to the pretrained policy. For example, Advantage Weighted Actor-Critic (AWAC) (Nair et al., 2020) and Implicit Q-Learning (IQL) (Kostrikov et al., 2021) applied offline regularization techniques directly to online fine-tuning. Building on the similar idea, PROTO (Li et al., 2023) introduced KL regularization to explicitly constrain the online policy to the pretrained one. Another line of work proposes new replay strategies for incorporating offline data more effectively. Lee et al. (2022) propose balanced replay to mitigate distribution shift and bootstrap error when transitioning from offline to online learning, Liu et al. (2024) employs a diffusion model to select or generate samples. Alternative strategies separate the roles of exploration and exploitation during fine-tuning. Jump-Start RL (JSRL) (Uchendu et al., 2023), maintains a fixed guided policy from pretraining alongside an exploration policy that is updated online, progressively transferring control from the pretrained policy to the learned one. More recent directions include expanding the action space via policy set expansion (PEX (Zhang et al., 2023)) and Bayesian methods for uncertainty-aware exploration (BOORL (Hu et al., 2024)). Some approaches skip offline pretraining but still make use of offline data. For example, Song et al. (2022) and Ball et al. (2023) start training directly with online RL while incorporating offline datasets, providing another way to combine offline data with online interaction.
Plasticity and Stability.
Plasticity and stability have long been recognized as central, often competing, objectives in learning systems. In neuroscience, this tension is formalized as the plasticity–stability dilemma (Mermillod et al., 2013), highlighting the challenge of integrating new knowledge without overwriting previously acquired competencies. In machine learning, similar dynamics manifest when agents must adapt to new data while preserving useful prior knowledge. Early work on continual and lifelong learning addressed this challenge via regularization techniques (Kirkpatrick et al., 2017), replay buffers (Rolnick et al., 2019) and modular architectures (Rusu et al., 2016). More recently, researchers have observed that insufficient plasticity can also hinder online deep RL, motivating methods designed to enhance plasticity of the neural network during training (Nikishin et al., 2022; Sokar et al., 2023; Dohare et al., 2024). In our work, we extend this perspective by framing offline-to-online RL explicitly as a stability–plasticity problem, where the central challenge is to balance the preservation of prior knowledge from pretraining with the adaptability needed to learn from new online experiences.
Appendix B Detailed Experimental Setup and Complete Results
B.1 offline pretraining
For offline pretraining, we train CalQL for 1M gradient steps on AntMaze, 20k on Adroit, and 250k on both Kitchen and MuJoCo locomotion tasks. ReBRAC is trained for 1M gradient steps on AntMaze, 100k on Adroit, 250k on Kitchen, and 500k on MuJoCo tasks. For the behavior cloning (BC) baseline, we perform 500K gradient steps of policy learning followed by 100k steps of fitted Q evaluation (Voloshin et al., 2019) (FQE) to obtain a Q-function for subsequent RL fine-tuning. Since this work primarily focuses on the online fine-tuning stage of offline-to-online RL, we do not modify the offline pretraining algorithm or its default hyperparameters.
For regime classification, we employ the two one-sided $t$-test (TOST) procedure with a margin of $\delta=0.05$ and a significance level of $\alpha=0.05$. The goal is to formally assess whether the pretrained policy $\pi_{0}$ and the offline dataset $\pi_{\mathcal{D}}$ are statistically indistinguishable in performance, or whether one is significantly superior. Let $\mu_{0}$ and $\mu_{\mathcal{D}}$ denote the mean returns of $\pi_{0}$ and $\pi_{\mathcal{D}}$. We conduct two one-sided tests for the null hypotheses $H_{0}:\mu_{0}-\mu_{\mathcal{D}}\leq-\delta$ and $H_{0}:\mu_{0}-\mu_{\mathcal{D}}\geq\delta$. If both null hypotheses are rejected, the difference is within the margin and the two are considered comparable, leading to assignment to the Comparable Regime. If only one hypothesis is rejected, the difference is statistically significant and exceeds the margin, and the setting is assigned to either the Superior or Inferior Regime depending on which policy achieves the higher mean return. The statistics for each dataset and pretraining policy are reported in Table 4.
While we use the margin parameter $\delta=0.05$ in the main experiments, we further assess the sensitivity to $\delta$ by comparing results under $\delta=0$ and $\delta=0.1$. The corresponding fine-tuning confusion matrices are reported in Table 2 and Table 3.
Table 2: Confusion matrix of fine-tuning results across the three pretraining regimes with margin $\delta=0$.
Pretraining Regime
Superior Comparable Inferior
Fine-tune
$\pi_{0}$-centric >> $\mathcal{D}$-centric 26 0 1
$\pi_{0}$-centric $\approx$ $\mathcal{D}$-centric 8 0 3
$\pi_{0}$-centric << $\mathcal{D}$-centric 4 1 20
Table 3: Confusion matrix of fine-tuning results across the three pretraining regimes with margin $\delta=0.1$.
Pretraining Regime
Superior Comparable Inferior
Fine-tune
$\pi_{0}$-centric >> $\mathcal{D}$-centric 18 9 0
$\pi_{0}$-centric $\approx$ $\mathcal{D}$-centric 4 5 2
$\pi_{0}$-centric << $\mathcal{D}$-centric 2 9 14
B.2 online fine-tuning
Across all environments, online fine-tuning is performed for 500k environment steps with UTD=1. The complete results, categorized according to the regime taxonomy, are reported in Figure 6, Figure 7, and Figure 8. To obtain the strongest performance for each class in each settings, we compare the interquartile mean (IQM) of evaluation results of online data warm-up and offline RL regularization within $\pi_{0}$-centric methods, and analogously compare offline data replay with and without reset within $\mathcal{D}$-centric methods. We then use the higher value from each class, comparing $\pi_{0}$-centric and $\mathcal{D}$-centric methods using two-sided $t$-tests with $\alpha=0.05$. To obtain stable and reliable $t$-test statistics, we base our analysis on the last 10 evaluation results from each random seed during online fine-tuning, which correspond to the final 50k training steps given our evaluation frequency of every 5k steps. An exception is made for door-binary-v0 and pen-binary-v0, where we instead use results up to 200k steps, since by the end of training nearly all methods achieve a 100% success rate, leaving no differences.
Figure 6: Full fine-tuning results in the Superior regime.
Figure 7: Full fine-tuning results in the Inferior regime.
Figure 8: Full fine-tuning results in the Comparable regime.
B.3 Compute Details
All experiments were conducted on a single-GPU setup using an NVIDIA L40S GPU, 24 CPU workers, and 20GB of RAM.
Table 4: Statistics (mean $\pm$ std) of offline datasets and pretrained policies. Pretrained policy scores $J(\pi_{0})$ are reported as averages over 10 random seeds.
Dataset Dataset CalQL ReBRAC BC
$J(\pi_{\mathcal{D}})$ #Trajs $J(\pi_{0})$ order $J(\pi_{0})$ order $J(\pi_{0})$ order
halfcheetah-random-v2 -0.001 $\pm$ 0.006 1000 0.248 $\pm$ 0.014 >> 0.275 $\pm$ 0.011 >> 0.019 $\pm$ 0.001 $\approx$
halfcheetah-medium-replay-v2 0.271 $\pm$ 0.135 202 0.451 $\pm$ 0.002 >> 0.504 $\pm$ 0.003 >> 0.370 $\pm$ 0.007 >>
halfcheetah-medium-v2 0.406 $\pm$ 0.029 1000 0.470 $\pm$ 0.003 >> 0.651 $\pm$ 0.010 >> 0.427 $\pm$ 0.002 $\approx$
halfcheetah-medium-expert-v2 0.643 $\pm$ 0.239 2000 0.519 $\pm$ 0.078 << 1.010 $\pm$ 0.019 >> 0.563 $\pm$ 0.025 <<
hopper-random-v2 0.012 $\pm$ 0.005 45240 0.091 $\pm$ 0.017 >> 0.082 $\pm$ 0.036 $\approx$ 0.038 $\pm$ 0.022 $\approx$
hopper-medium-replay-v2 0.150 $\pm$ 0.157 2039 1.001 $\pm$ 0.009 >> 0.965 $\pm$ 0.042 >> 0.398 $\pm$ 0.037 >>
hopper-medium-v2 0.443 $\pm$ 0.117 2187 0.672 $\pm$ 0.039 >> 1.016 $\pm$ 0.012 >> 0.554 $\pm$ 0.010 >>
hopper-medium-expert-v2 0.648 $\pm$ 0.319 3214 1.058 $\pm$ 0.108 >> 1.063 $\pm$ 0.042 >> 0.556 $\pm$ 0.012 <<
walker2d-random-v2 0.000 $\pm$ 0.001 48908 0.082 $\pm$ 0.043 >> 0.058 $\pm$ 0.001 >> 0.008 $\pm$ 0.001 $\approx$
walker2d-medium-replay-v2 0.148 $\pm$ 0.195 1093 0.843 $\pm$ 0.025 >> 0.853 $\pm$ 0.045 >> 0.276 $\pm$ 0.074 >>
walker2d-medium-v2 0.620 $\pm$ 0.239 1191 0.742 $\pm$ 0.071 >> 0.845 $\pm$ 0.008 >> 0.507 $\pm$ 0.073 <<
walker2d-medium-expert-v2 0.826 $\pm$ 0.285 2191 1.073 $\pm$ 0.035 >> 1.113 $\pm$ 0.004 >> 1.075 $\pm$ 0.004 >>
pen-binary-v0 1.000 $\pm$ 0.000 846 0.657 $\pm$ 0.059 << 0.451 $\pm$ 0.086 << 0.589 $\pm$ 0.068 <<
door-binary-v0 1.000 $\pm$ 0.000 82 0.112 $\pm$ 0.123 << 0.000 $\pm$ 0.000 << 0.000 $\pm$ 0.000 <<
relocate-binary-v0 1.000 $\pm$ 0.000 36 0.010 $\pm$ 0.011 << 0.000 $\pm$ 0.000 << 0.000 $\pm$ 0.000 <<
kitchen-partial-v0 0.586 $\pm$ 0.187 600 0.764 $\pm$ 0.094 >> 0.133 $\pm$ 0.085 << 0.222 $\pm$ 0.079 <<
kitchen-mixed-v0 0.598 $\pm$ 0.145 600 0.464 $\pm$ 0.092 << 0.034 $\pm$ 0.033 << 0.275 $\pm$ 0.049 <<
kitchen-complete-v0 1.000 $\pm$ 0.000 19 0.043 $\pm$ 0.078 << 0.002 $\pm$ 0.004 << 0.385 $\pm$ 0.202 <<
antmaze-large-diverse-v2 0.106 $\pm$ 0.308 999 0.305 $\pm$ 0.055 >> 0.399 $\pm$ 0.080 >> 0.000 $\pm$ 0.000 <<
antmaze-large-play-v2 0.105 $\pm$ 0.307 999 0.247 $\pm$ 0.068 >> 0.351 $\pm$ 0.072 >> 0.000 $\pm$ 0.000 <<
antmaze-ultra-diverse-v2 0.053 $\pm$ 0.224 999 0.118 $\pm$ 0.072 $\approx$ 0.127 $\pm$ 0.132 $\approx$ 0.000 $\pm$ 0.000 $\approx$